QUESTION: A study believes that 70% of adults in the Philippines own a cellphone.

A cellphone manufacturer believes that the actual number is much

less than 70%. 100 Filipino adults were surveyed, of which 74 have

cellphones. Using a 5% level of significance, is the cellphone

manufacturer’s claim valid or not?


CLAIM:

EVIDENCE:

REASONING:


1
Expert's answer
2022-06-01T11:39:50-0400

Claim: the actual number is much less than 70%.

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p\ge0.70"

"H_a:p<0.70"

This corresponds to a left-tailed test, for which a z-test for one population proportion will be used.

Evidence:

Based on the information provided, the significance level is "\alpha = 0.05

," and the critical value for a left-tailed test is "z_c = -1.6449."

The rejection region for this left-tailed test is "R = \{z: z < -1.6449\}."

The z-statistic is computed as follows:



"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{74/100-0.7}{\sqrt{\dfrac{0.7(1-0.7)}{100}}}=2.1822"

Since it is observed that "z = 2.1822 \ge-1.6449= z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p=P(Z<2.1822)= 0.985453," and since "p= 0.985453>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Reasoning:

Therefore, there is not enough evidence to claim that the population proportion "p" is less than 0.70, at the "\alpha = 0.05" significance level.